HUBBLE SPACE TELESCOPE DAILY REPORT #1256 PERIOD COVERED: 0000Z Daily Status Report as of 305/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5309 (Star Formation and the Massive Star Population in Blue Compact Dwarf Galaxies) The WF/PC-2 was used to make multicolor images of the blue compact dwarf (BCD) galaxy IZW18 (aka, One Zwicky 18) which is known to have a low abundance of heavy elements. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Four Sets of FOS/BL 5444 (A UV Spectral Library of Metal- Poor Massive Stars) The FOS (Blue Side) is currently operating on low voltage. No science observations were taken. 1.3 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of NGC-1019 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOC/48 5763 (FOC/48 Long Slit Aperture Location) The Objective of this CAL/FOC proposal is to measure the location of the FOC f/48 long slit aperture during the approximate three hour period in which the instrument is usable after turn on. Cluster NGC-1068 was viewed. The observations were completed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5458 (Evolution of Galaxies in Intermediate Redshift Clusters) The WF/PC-2 was used to obtain deep exposures of galactic cluster ABELL0665 to measure its dynamical parameters. the observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Notes: An anomaly-free operational day was experienced by each HST subsystem. The FOS (Blue Side) high voltage is scheduled to be turned-on Saturday, Nov. 5, 1994 at 1330Z.. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-